OpenStudy (anonymous):

(2)/(x-4)(x-2) what is the X intercept?

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

is the equation \(\huge \rm y=\frac{2}{(x-4)(x-2)}\)

OpenStudy (anonymous):

F(x)=

OpenStudy (anonymous):

there aren't any x-intercepts

OpenStudy (anonymous):

Why not

OpenStudy (anonymous):

yes no x-intercepts to get x intercepts means to solve f(x)=0 if we set \(\huge \rm \frac{2}{(x-2)(x-4)}=0\) we can only have the top equal to zero since zero at the bottom will make our function undefined. However, in this case, the top is 2 which different than zero so no x-intercepts

OpenStudy (anonymous):

Are there any y interceptors?

OpenStudy (anonymous):

x=0 the y= 1/4 in fact the function never touches zero

OpenStudy (anonymous):

yes there is that is the one i just wrote x=0 y=1/4 is y intercept

OpenStudy (anonymous):

I take that back, you're right Eurler I thought it said 2x in the numerator

OpenStudy (anonymous):

No x - intercepts

OpenStudy (anonymous):

Yeah it was 2 only! i know you get it confused lol

OpenStudy (anonymous):

=P

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Good job Euler ^.^

OpenStudy (anonymous):

^.^

OpenStudy (anonymous):

Essentially when they ask for intercepts, if they say they want x, set y = 0 and solve for x. If they want y, set x = 0 and solve for y.

OpenStudy (anonymous):

I can confirm Euler on there being no x - intercepts and y being 1/4.